Find the asymptotes of y^2(a+x) =x^2(3a-x) parallel to coordinate axes



1
Expert's answer
2022-03-23T06:43:46-0400

"Horyzontal\\,\\,asymptotes:\\\\\\underset{x\\rightarrow \\infty}{\\lim}y\\left( x \\right) =C\\\\C^2=\\underset{x\\rightarrow \\infty}{\\lim}\\frac{x^2\\left( 3a-x \\right)}{a+x}=\\infty \\\\No\\,\\,horyzontal\\,\\,asymptotes\\\\Vertical\\,\\,asymptotes:\\\\\\underset{x\\rightarrow x_0}{\\lim}y\\left( x \\right) =\\infty \\Rightarrow \\underset{x\\rightarrow x_0}{\\lim}y^2\\left( x \\right) =+\\infty \\Rightarrow \\\\\\Rightarrow \\underset{x\\rightarrow x_0}{\\lim}\\frac{x^2\\left( 3a-x \\right)}{a+x}=+\\infty \\Rightarrow \\left\\{ \\begin{array}{c}\ta+x=0\\\\\t3a-x\\ne 0\\\\\\end{array} \\right. \\\\Thus\\,\\,for\\,\\,a=0 no\\,\\,vertical\\,\\,asymptotes\\\\For\\,\\,a\\ne 0: x=-a\\,\\,-vertical\\,\\,asymptote"